1)A profit maximizing Monopolist has the following total revenue and total cost function

TR=100q-2q²

TC=50+40q

Where q=output level

You are required to

1.derive the marginal cost

2.derive the marginal revenue

3.calculate the profit maximizing output level

4.determine the profit maximizing selling price.


1
Expert's answer
2021-07-23T15:43:02-0400

MC"=" "\frac{d}{dQ}" "(" TC")" "=" "\frac{d}{dQ}" "(" 50"+" 40q")" "=" 40

MR"=""\frac{d}{dQ}""(" TR")" "=" "\frac{d}{dQ}" "(" 100q"-" 2q"^2" ")" "="100"-" 4q

Firm maximize profit when MR"="MC

100"-" 4q"=" 40

q"=" 15 units


Profit maximizing price:

100"-" 4q"=" 100"-" 4"(" 15")" "=" 40